GATE प्रश्न और उत्तर का अभ्यास करें

प्र: Rajitha's present age is three years less than twice her age 12 years ago. Also, the ratio of Raj's present age to Rajitha's present age is 4:9. What will be Raj's age after 5 years ? 2341 0

  • 1
    17 yrs
    सही
    गलत
  • 2
    16 yrs
    सही
    गलत
  • 3
    19 yrs
    सही
    गलत
  • 4
    21 yrs
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "17 yrs"
व्याख्या :

Answer: A) 17 yrs Explanation: Let the present age of Raj and Rajitha be 4x and 9x.Then, according to question,9x = 2(9x-12)-3 => 9x = 27 => x = 3Raj's present age = 4x = 12yrs, After 5 years Raj's age will be 12 + 5 = 17 yrs.

प्र: Which revolutionary died after a 64 days fast in jail? 2340 0

  • 1
    Jatin Das
    सही
    गलत
  • 2
    Pulin Das
    सही
    गलत
  • 3
    Subhash Chandra bose
    सही
    गलत
  • 4
    Both A & B
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "Jatin Das"
व्याख्या :

Answer: A) Jatin Das Explanation: Jatindra Nath Das, also known as Jatin Das, was a revolutionary and Independence activist. He embraced death in Lahore jail after fasting for 64 days in 1929 at the age of 24 yrs.

प्र: 1, 2, 8, 33, ? Find the next number in the given series? 2340 0

  • 1
    148
    सही
    गलत
  • 2
    145
    सही
    गलत
  • 3
    143
    सही
    गलत
  • 4
    141
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "148"
व्याख्या :

Answer: A) 148 Explanation: 1 x 1 + 1 x 1 = (1 + 1) = 2 2 x 2 + 2 x 2 = (4 + 4) = 8 8 x 3 + 3 x 3 = (24 + 9) = 33 33 x 4 + 4 x 4 = (132 + 16) = 148 148 x 5 + 5 x 5 = (740 + 25) = 765.

प्र: Derivative of sinx by first principle? 2339 0

  • 1
    -sinx
    सही
    गलत
  • 2
    -cosx
    सही
    गलत
  • 3
    cosx
    सही
    गलत
  • 4
    sinx
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 3. "cosx"
व्याख्या :

Answer: C) cosx Explanation:   Thus, the derivative of sinx by first principle is cosx.

प्र: Which statement is one component of the cell theory? 2339 0

  • 1
    DNA is the basis of genetic transfer
    सही
    गलत
  • 2
    All living cells arise from pre-existing cells
    सही
    गलत
  • 3
    All cells are composed of same organelles
    सही
    गलत
  • 4
    Tissues are formed from combinations of cells
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 2. "All living cells arise from pre-existing cells"
व्याख्या :

Answer: B) All living cells arise from pre-existing cells Explanation: All living cells arise from pre-existing cells is one component of the cell theory.   The Cell Theory is one of the basic principles of biology.   The Cell Theory states: 1. All living organisms are composed of cells. They may be unicellular or multicellular. 2. The cell is the basic unit of life. 3. Cells arise from pre-existing cells.

प्र: A problem is given to three students whose chances of solving it are 1/2, 1/3 and 1/4 respectively. What is the probability that the problem will be solved? 2338 0

  • 1
    1/4
    सही
    गलत
  • 2
    1/2
    सही
    गलत
  • 3
    3/4
    सही
    गलत
  • 4
    7/12
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 3. "3/4"
व्याख्या :

Answer: C) 3/4 Explanation: Let A, B, C be the respective events of solving the problem and A , B, C be the respective events of not solving the problem. Then A, B, C are independent event ∴A, B, C are independent events Now,  P(A) = 1/2 , P(B) = 1/3 and P(C)=1/4  PA=12, PB=23, PC= 34 ∴ P( none  solves the problem) = P(not A) and (not B) and (not C)                     = PA∩B∩C                    = PAPBPC         ∵ A, B, C are Independent                                          =  12×23×34                     = 14   Hence, P(the problem will be solved) = 1 - P(none solves the problem)                  = 1-14= 3/4

प्र: In a group of 6 boys and 4 girls, four children are to be selected. In how many different ways can they be selected such that at least one boy should be there ? 2338 0

  • 1
    205
    सही
    गलत
  • 2
    194
    सही
    गलत
  • 3
    209
    सही
    गलत
  • 4
    159
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 3. "209"
व्याख्या :

Answer: C) 209 Explanation: We may have (1 boy and 3 girls)or(2boys and 2 girls)or(3 boys and 1 girl)or(4 boys). Required number of ways = (C16× C34) + C26×C24  + (C36× C14) + (C46)   = (24+90+80+15)  = 209.

प्र: Find the odd one in the given series ? 5-20-25, 10-40-50, 4-15-24, 25-100-125 2338 0

  • 1
    4-15-24
    सही
    गलत
  • 2
    5-20-25
    सही
    गलत
  • 3
    10-40-50
    सही
    गलत
  • 4
    25-100-125
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "4-15-24"
व्याख्या :

Answer: A) 4-15-24 Explanation: In each term second two numbers are multiples of the first number. But in 4-15-24 15 is not multiple of 4.

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई